OpenStudy (anonymous):

Explain how to graph x – 2y = 4 using the x- and y-intercepts method. Be sure to use complete sentences and use the x- and y-intercepts method only.

OpenStudy (anonymous):

Okay. We have our equation here: x - 2y = 4 ~~~~~~~~~ There are two rules we need to know: 1. Find the X intercept by making Y = 0. 2. Find the Y intercept by making X = 0. ~~~~~~~~~~~ So, here's how we find the X intercept: x - 2(0) = 4 See how I made Y equal 0? So, x - 0 = 4 X = 4. So, in this case, the x intercepts would be..... (4, 0) and (-4, 0) ~~~~~~~~~~~~~~ You can use the second rule to find the Y intercept. Just make X equal to 0.

OpenStudy (anonymous):

But how would I explain that?

OpenStudy (anonymous):

Sorry for having left earlier. My tutoring client arrived, and I needed to get to work. You could explain it the same way I did, above. In order to find the Y intercept, the first thing you would need to do is make X equal to 0. When you finally do have the value for the Y intercept, remember to put it in the Y slot, in the ordered pairs. For example, if the Y intercept were 4... then these would be the Y intercept ordered pairs: (0, 4) and (0, -4).